If a random variable X follows Normal distribution such that P (9.65<=X<=13.8)=0.07008and P(X>9.6)=0.8159 where Find the mean & variance of X and find P(X>=13.8 /X>=9.6).
1
Expert's answer
2020-12-29T15:10:44-0500

"P(X>9.6) = P(Z > \frac {9.6 - \mu} {\sigma}) =\\
1 - P(Z \le \frac {9.6 - \mu} {\sigma}) = 0.8159\\
\frac {9.6 - \mu} {\sigma}=Z_{1-0.8159}=Z_{0.1841} =-0.8999"



"P(9.6\le X\le13.8) = P(X\ge 9.6) - P(X\ge 13.8) = 0.07008\\
P(X \ge 13.8) = 0.8159 - 0.07008 = 0.74582\\
\frac {13.8 - \mu} {\sigma}=Z_{1-0.74582}=Z_{0.25418} =-0.6614"


From two equations we found that sigma = 17.6133, mean = mu = 25.4493,

var = s2=310.227


And

"P(X\ge13.8 |X\ge 9.6) =\frac{P(X\ge13.8)}{P(X\ge 9.6)}=\frac{0.74582}{0.8159}= 0.9141"